# Time-resolved magneto-optical Kerr effect in the altermagnet candidate MnTe

### v1

@@ Abstract
$\alpha$-MnTe is an antiferromagnetic semiconductor with above room temperature $T_{N}$ = 310 K, which is promising for spintronic applications. Recently, it was predicted to be an altermagnet, containing bands with momentum-dependent spin splitting; time-resolved experimental probes of magnetism in MnTe are therefore important both for understanding the magnetic structure and potential device applications. We investigate ultrafast spin dynamics in epitaxial MnTe(001)/InP(111) thin films using the time-resolved magneto-optical Kerr effect. At room temperature, we observe an oscillation mode at 55 GHz that does not appear at zero magnetic field. Combining field, polarization, and temperature dependence, we identify this mode as an acoustic phonon-coupled magnon, likely originating from inverse stimulated Raman scattering. Additionally, we observe two optical phonons at 3.6 THz and 4.2 THz, which broaden and redshift with increasing temperature.
@@
In the past several years, antiferromagnets have been developed into active elements of spintronic devices, forming memory elements Wadley et al. (2016); Higo et al. (2022), terahertz nano-oscillators Cheng, Xiao, and Brataas (2016); Olejník et al. (2018), and low-dissipation interconnects in hybrid quantum systems Lebrun et al. (2018). A natural route towards integrating these devices with traditional semiconductor functionality is to employ magnetic semiconductors Jungwirth et al. (2016); while ferromagnetic semiconductors that are ordered at room temperature are rare, antiferromagnetic semiconductors are more common and offer new methods of controlling the spin degree of freedom, such as coupling between magnons and excitons Bae et al. (2022). An example material is hexagonal $\alpha$-MnTe, promising due to its high Néel temperature $T_{N}$approx 310 K Szuszkiewicz et al. (2005), stable magnetoresistance reflecting Néel orientation Kriegner et al. (2016), and spin-phonon coupling Zhang et al. (2020a); Bossini et al. (2021). Very recently, MnTe has received increased attention after the prediction that it is not simply an antiferromagnet, but an altermagnet Šmejkal, Sinova, and Jungwirth (2022a, b). In this new magnetic phase, the spin-up and spin-down bands are neither degenerate, like conventional antiferromagnets, nor uniformly split, like ferromagnets; rather, the bands contain anisotropic spin polarization that alternates sign in momentum space while preserving zero net spin polarization.
[…]
@@
In this letter, we perform time-resolved magneto-optical Kerr effect (TR-MOKE) measurements on hexagonal MnTe thin films as a function of temperature, polarization, and magnetic field. We identify optical phonon modes at 3.6 THz and 4.2 THz and show that they broaden and redshift with increasing temperature from 10 K to 350 K. In the presence of magnetic field, we observe an oscillation at 55 GHz which we hypothesize is due to an acoustic phonon-coupled magnon. This oscillation can be enhanced or suppressed by changing the pump polarization angle, which indicates a non-thermal origin. Oscillation components at 55 GHz persist even above the nominal $T_{N}$ = 310 K, up to the maximum measured $T$ = 335 K, which indicates that acoustic phonons as well as the magnon contribute to the TR-MOKE signal.
@@
The structure of hexagonal $\alpha$-MnTe, with space group $P6_{3}/mmc$, is shown in Fig. 1(a). It is an A-type antiferromagnet, containing ferromagnetic planes of in-plane Mn spins that alternate spin direction along the $c$-axis. The easy axis is at ${}^{\circ}$ to the crystalline $a$-axis. Our samples are 40 nm-thick MnTe(001) thin films, grown on InP(111) via molecular beam epitaxy.
@@
We perform two-color pump-probe TR-MOKE, schematically illustrated in Fig. 1(b). We use a 1560 nm pump at 10 mW power and a 780 nm probe at 0.7 mW power on the sample; both pump and probe pulses are $\approx$100 fs in duration and have 80 MHz repetition rate. Both beams are focused through a 10x microscope objective at normal incidence, forming a 30 $\mu$m spot size. The reflected light is short-pass filtered to reject the pump, then the polarization rotation $\Delta\theta_{K}(t)$ is measured with a Wollaston prism and a balanced photodiode. The signal-to-noise is enhanced by modulating the pump between orthogonal polarization states with a PEM at 42 kHz (placed before the half-wave plate that controls pump polarization) and measuring the second harmonic of the photodiode voltage output using a lock-in amplifier. Both pump and probe are linearly polarized.
@@
Figure 1: Time-resolved magneto-optical Kerr effect (TR-MOKE) setup and observation of optical phonons in MnTe. (a) Crystal and magnetic structure of hexagonal MnTe (001). Gold and silver balls represent Mn and Te, respectively. (b) Schematic of two-color pump-probe MOKE with 1560 nm pump and 780 nm probe, both $\sim$100 fs in duration. (c) TR-MOKE scans of MnTe(001) at 295 K as a function of linear pump polarization, and (d) associated fast Fourier transforms (FFTs).
@@
We first show short time scans (up to 13 ps) at room temperature (295 K) as a function of pump polarization angle in Fig. 1(c). After an initial sharp peak, we observe beating between two modes. Upon taking the FFTs in Fig. 1(d), we find peaks at 3.6 and 4.2 THz. These frequencies match modes observed in previous Raman spectroscopy of MnTe Zhang et al. (2020a), where they were attributed to transverse and longitudinal optical phonons, respectively. We find that varying the pump polarization changes the relative amplitude of each mode. Fitting the oscillations to two damped sines, we find that at least one mode shows three-fold symmetry (shown in the supporting information), which one might expect from the hexagonal lattice. Polarization-dependent amplitudes demonstrate a non-thermal excitation mechanism, such as impulsive stimulated Raman scattering (ISRS) Yan, Gamble, and Nelson (1985) or displacive excitation of coherent phonons (DECP) Zeiger et al. (1992). Since DECP requires absorption of a photon (meaning above-bandgap excitation), ISRS does not, and our pump wavelength is below the bandgap of MnTe(1.15 eV = 1078 nm at 295 K Bossini et al. (2020)), ISRS is the most likely excitation mechanism. Our results differ from previous time-resolved reflectivity and MOKE on MnTe Bossini et al. (2021); Deltenre et al. (2021), where only the doubly degenerate $E_{2g}$ phonon mode at 5.3 THz (also seen in Raman scattering Zhang et al. (2020a)) was observed. In these works, however, the pump wavelength was 800 nm (above the bandgap), the probe wavelength was 730 nm, and the pump-probe signal was explicitly attributed to DECP. These different excitation mechanisms and probe wavelengths might cause the observed differences in phonon spectra.
[…]
@@
Figure 2: Magnon and acoustic phonon oscillations at 295 K. (a) Longer scans (133 ps duration) at fixed ${}^{\circ}$ pump angle as a function of out-of-plane magnetic field $H_{app}$, and (b) associated FFTs. We observe a mode at 55 GHz with field-dependent amplitude, which indicates spin oscillations, as well as other features in the FFT at $\sim$17 GHz and 34 GHz that may represent acoustic phonons. (c) Pump polarization dependence of $\Delta\theta_{K}(t)$ at fixed $H_{app}=2$ kG, with FFTs in (d). The oscillations are suppressed or enhanced as a function of pump angle, demonstrating a non-thermal origin such as phonon-induced magnons. (e) Difference and sum of $\Delta\theta_{K}(t)$ at +2.2 kG and -2.2 kG, and (f) associated FFTs. Both plots contain oscillation components at 55 GHz, which means non-magnetic signal cannot be straightforwardly subtracted out by subtracting opposite fields.
@@
We fit the data from 10 ps onwards, discarding the initial phonon modes, to the sum of an exponential decay representing electron-phonon relaxation Bossini et al. (2021), two or three damped oscillations (depending on the pump angle, see the supporting information) that represent acoustic phonons and/or magnons Gamaly and Rode (2013), and a linear background:
$\begin{split}\Delta\theta_{K}(t)=A_{exp}e^{-t/t_{0}}+A_{1}e^{-\alpha_{1}t}\sin\left(2\pi f_{1}t+\phi_{1}\right)&\\ +A_{2}e^{-\alpha_{2}t}\sin\left(2\pi f_{2}t+\phi_{2}\right)&\\ +A_{3}e^{-\alpha_{3}t}\sin\left(2\pi f_{3}t+\phi_{3}\right)+Bt+C\end{split}$ (1)
@@
Fits are shown with the data in Fig. 2(a), except at $H_{app}$ = 0 kG, where a consistent fit could not be obtained. At this pump angle (${}^{\circ}$), the features of the data are largely captured by two damped oscillations and using three did not significantly improve the fit quality (see the supporting information). The linear background varies from run to run and is attributed to experimental artifacts, for example from inevitable slight defocusing of the probe beam when scanning the delay line. Best fit parameters are shown in Table 1. We find that although both oscillations go to zero when $H_{app}=0$, $A_{1}$ monotonically depends on the magnitude of $H_{app}$, which suggests a magnetic origin, whereas $A_{2}$ does not.
Previous studies of magnons in MnTe are scarce; there is one neutron scattering report of a zone boundary magnon at 5.6 THz at 295 K Szuszkiewicz et al. (2005), while A DFT study Mu et al. (2019) calculates $\approx$ 0.5 THz at $k=0$. We can estimate the $k=0$ magnon frequencies from the anisotropy, which in MnTe is largely easy-plane. Within the easy plane there are nominally three equivalent easy axes; however, MnTe on InP(111) experiences biaxial tensile strain due to the difference in thermal coefficient between InP and MnTe Kriegner et al. (2017). This strain modifies the in-plane anisotropy in a not-well-understood manner that depends also on sample thickness. Regardless, since the in-plane anisotropy field $H_{Az}$ is small compared to the hard-axis (easy-plane) field $H_{Ax}$, we can use the formula for the two $k=0$ easy-plane magnon frequencies $\omega_{\alpha}$ and $\omega_{\beta}$ Rezende, Azevedo, and Rodríguez-Suárez (2019):
@@
where $\gamma=g\mu_{B}/\hbar$ is the gyromagnetic ratio, $H_{E}$ is the exchange field, and $H_{Ax}$ and $H_{Az}$ are the hard-axis and easy-axis anisotropy fields, respectively. For $H_{Ax}>>H_{Az}$, Eqn. 2 reduces to:
@@
Although the values of $H_{Az}$ and $H_{Ax}$ are not well-known, we can get an estimate by using the spin-flop field $H_{SF}\approx\sqrt{2H_{E}H_{Az}}$, which was measured to be 2 T for MnTe/InP(111) Kriegner et al. (2017). Assuming $g=2$ for MnTe, the in-plane magnon frequency $\omega_{\beta 0}$ evaluates to 56 GHz, consistent with our measured value. From the lack of any measurable magnetization response in MnTe with OOP field up to 6 T in Ref. Kriegner et al., 2016, we can estimate that $H_{Ax}>6$ T, which means that the out-of-plane magnon frequency $\omega_{\alpha 0}>$ 176 GHz - therefore, features between 17 and 35 GHz are unlikely to be magnons. They may represent acoustic phonons, or a non-monotonic magneto-optical response to sample heating Yang et al. (2019). In previous studies of the easy-plane antiferromagnet NiO, the in-plane mode was observed with a linearly polarized probe through Faraday rotation, while the out-of-plane mode was observed with a circularly polarized probe through the Cotton-Mouton effect (linear birefringence) Tzschaschel et al. (2017). If similar mechanisms are active in MnTe, the higher-frequency magnon may only appear when using a circularly polarized probe, which will be the subject of future work.
@@
Following the field dependence, we measure pump polarization dependence of the spin oscillations at fixed $H_{app}$ = 2.2 kG in Fig. 2(c), with associated FFTs in 2(d). We find that the oscillation amplitudes are pump polarization-dependent - strongest at ${}^{\circ}$ and ${}^{\circ}$, and nearly disappearing at ${}^{\circ}$ and ${}^{\circ}$ (although at the current signal-to-noise level, we cannot definitively assign them three-fold symmetry). We also note that data at some angles, such as ${}^{\circ}$, fit better to three oscillations, while data at other angles, such as ${}^{\circ}$ and ${}^{\circ}$, only seem to contain two. This suggests that multiple acoustic phonons with different pump angle dependences may be excited by the laser.
@@
Pump polarization dependence of the GHz oscillations establishes a non-thermal origin, as with the optical phonons. In general, magnons measured with TR-MOKE can be thermally or non-thermally generated. In thermal generation, heating from the pump laser temporarily changes the anisotropy field $H_{A}$, which reorients the effective field $H_{eff}$ and causes the spins to precess around the new $H_{eff}$ Lyalin, Cheng, and Kawakami (2021). Non-thermal magnon generation is often caused by spin-phonon coupling in magnetic semiconductors and insulators when the laser coherently excites phonons Berk et al. (2019), and were observed in MnTe in Refs. Bossini et al., 2021; Deltenre et al., 2021. In those experiments, the laser excited the 5.3 THz $E_{2g}$ optical phonon, causing the Te atoms to oscillate out-of-phase. This modulated the Mn-Te-Mn bond angle, in turn causing an oscillation in the weak Te-mediated superexchange $J_{3}$ which generated magnons at the same frequency. In our data, however, the 3.6 THz and 4.2 THz oscillations in $\Delta\theta_{K}(t)$ show no field dependence within experimental resolution (see the supplemental information). Therefore, we attribute them entirely to optical phonons. The polarization dependence in Fig. 2(c) is consistent with a hypothesis of acoustic phonon-induced spin oscillations at 55 GHz, driven by ISRS Yan, Gamble, and Nelson (1985).
@@
In some spin-phonon coupled systems, $\Delta\theta_{K}(t)$ contains contributions at $H_{app}=0$ from the phonons themselves (as in Fig. 1) or from phonon-induced spin tilting Lattery (2020); Zhang et al. (2020b). Typically, magnetic and non-magnetic signals are separated by taking the sum and difference of data acquired at opposite magnetic fields McGill et al. (2004). In Fig. 2(e), we plot the sum and difference of $\Delta\theta_{K}(t)$ at $H_{app}=+2.2$ kG and -2.2 kG, with FFTs in 2(f). We find oscillation components at 55 GHz in both quantities, which is consistent with a non-magnetic acoustic phonon and a coupled magnon at the same frequency that both contribute to the $\Delta\theta_{K}(t)$ signal. Therefore, isolating magnetic from non-magnetic contributions to $\Delta\theta_{K}(t)$ cannot be straightforwardly done by taking the difference and sum at low fields. Instead, we measure above $T_{N}$ ($\approx$ 300 K in our thin films Brahlek and Chilcote (2023), slightly lowered from bulk $T_{N}$ = 310 K due to strain), where we expect spin oscillations to disappear in a paramagnetic phase. In Fig. 3, we measure $\Delta\theta_{K}(t)$ time traces from 295 K to 335 K. We show scans at fixed $H_{app}$ = +2.2 kG from 290 335 K in 3(a), and FFTs in 3(b). Surprisingly, oscillation components at 55 GHz persist at least up to 335 K, the highest temperature measured. At this temperature, the oscillations are non-magnetic; yet, a clear field dependence was seen at room temperature in Fig. 2. Our hypothesis is that $\Delta\theta_{K}(t)$ is due to a phonon-coupled magnon below $T_{N}$, and purely the acoustic phonon - at the same frequency - above $T_{N}$. Increased contribution of non-magnetic phonons to $\Delta\theta_{K}(t)$ at 335 K compared to 295 K could be caused by several factors, such as changes in the phonon density of states or the birefringence coefficient at elevated temperature. Future work will investigate these oscillations further with high-temperature magnetic field dependence.
Figure 3: Coupled spin-phonon oscillations above room temperature. (a) $\Delta\theta_{K}(t)$ at fixed $H_{app}$ = +2.2 kG as a function of temperature, and (b) corresponding FFTs. Oscillations persist even above the nominal $T_{N}$ = 305-310 K, indicating phonon contributions to $\Delta\theta_{K}(t)$ above $T_{N}$.
@@
Lastly, we measure $\Delta\theta_{K}(t)$ as a function of temperature from 10 K up to 350 K in a cryostat. In this configuration, the maximum attainable $H_{app}$ was $\approx$ 500 G, which we found was not sufficient to resolve spin oscillations. Instead, we focus on the optical phonon oscillations, showing $\Delta\theta_{K}(t)$ traces in Fig. 4(a) and FFTs in 4(b). (Full time traces showing no spin oscillations are shown in the Supporting Information.)
@@
To analyze the temperature dependence of the phonons, we fit the background, then subtract it off and fit the residual oscillations to two damped sinusoids. We find that this two-step process is more robust than fitting the background and oscillations simultaneously. At low temperatures, modeling the background by a sum of decaying exponentials does not yield a satisfactory fit. Therefore, we adapt a phenomenological model developed to to fit pump-probe reflectivity in Ref. Djordjevic et al., 2006:

### v2

# Time-resolved magneto-optical effects in the altermagnet candidate MnTe

@@ Abstract
$\alpha$-MnTe is an antiferromagnetic semiconductor with above room temperature $T_{N}$ = 310 K, which is promising for spintronic applications. Recently, it was reported to be an altermagnet, containing bands with momentum-dependent spin splitting; time-resolved experimental probes of MnTe are therefore important both for understanding novel magnetic properties and potential device applications. We investigate ultrafast spin dynamics in epitaxial MnTe(001)/InP(111) thin films using pump-probe magneto-optical measurements in the Kerr configuration. At room temperature, we observe an oscillation mode at 55 GHz that does not appear at zero magnetic field. Combining field and polarization dependence, we identify this mode as a magnon, likely originating from inverse stimulated Raman scattering. Magnetic field-dependent oscillations persist up to at least 335 K, which could reflect coupling to known short-range magnetic order in MnTe above $T_{N}$. Additionally, we observe two optical phonons at 3.6 THz and 4.2 THz, which broaden and redshift with increasing temperature.
@@
In the past several years, antiferromagnets have been developed into active elements of spintronic devices, forming memory elements Wadley et al. (2016); Higo et al. (2022), terahertz nano-oscillators Cheng, Xiao, and Brataas (2016); Olejník et al. (2018), and low-dissipation interconnects in hybrid quantum systems Lebrun et al. (2018). A natural route towards integrating these devices with traditional semiconductor functionality is to employ magnetic semiconductors Jungwirth et al. (2016); while room-temperature ferromagnetic semiconductors are rare, antiferromagnetic semiconductors are more common and offer new methods of controlling the spin degree of freedom, such as coupling between magnons and excitons Bae et al. (2022). Hexagonal $\alpha$-MnTe is promising due to its high Néel temperature $T_{N}$approx 310 K Szuszkiewicz et al. (2005), stable magnetoresistance reflecting Néel orientation Kriegner et al. (2016), and spin-phonon coupling Zhang et al. (2020a); Bossini et al. (2021). Very recently, MnTe has received increased attention after the prediction Šmejkal, Sinova, and Jungwirth (2022a, b) and subsequent confirmation Krampaský et al. (2024) that it is not simply an antiferromagnet, but an altermagnet. In this new magnetic phase, the spin-up and spin-down bands are neither degenerate, like conventional antiferromagnets, nor uniformly split, like ferromagnets; rather, the bands contain anisotropic spin polarization that alternates sign in momentum space while preserving zero net spin polarization.
This anisotropic spin polarization has very recently been observed via ARPES in MnTe Krampaský et al. (2024), MnTe2 Zhu et al. (2024), and RuO2 Fedchenko et al. (2024), and is predicted to lead to new effects of both scientific and practical importance, such as an anomalous Hall effect Šmejkal, Sinova, and Jungwirth (2022a); Hayami and Kusunose (2021), an associated magneto-optical Kerr effect (MOKE), novel forms of spin-valley control Ma et al. (2021), and spin current generation Šmejkal et al. (2023); Hayami, Yanagi, and Kusunose (2019). Currently there are few experimental demonstrations of these effects, mostly in RuO2 Feng et al. (2022); Liu et al. (2023); Bai et al. (2022); Bose et al. (2022); Fedchenko et al. (2024). In MnTe, there is one report of an anomalous Hall effect attributed to altermagnetismGonzalez Betancourt et al. (2023); however, variations in stoichiometry and strain can lead to weak ferromagnetism, which can also produce an anomalous Hall effect Chilcote et al. . Meanwhile, MOKE in MnTe has been predicted Mazin (2023) but not yet observedChilcote et al. . Experimental studies of magnetism in MnTe are therefore important for better understanding of a potential altermagnetic phase.
@@
In this letter, we perform time-resolved magneto-optical (MO) measurements in the Kerr geometry on hexagonal MnTe thin films as a function of temperature, polarization, and magnetic field. We identify optical phonon modes at 3.6 THz and 4.2 THz and show that they broaden and redshift with increasing temperature from 10 K to 350 K. In the presence of magnetic field, we observe an oscillation mode at 55 GHz below $T_{N}$, which we hypothesize is due to an in-plane $k=0$ magnon, as the oscillation is absent at zero magnetic field Rezende, Azevedo, and Rodríguez-Suárez (2019). The oscillation amplitude depends on the pump polarization angle, which indicates a non-thermal origin. Surprisingly, magnetic field-dependent 55 GHz oscillations persist even above $T_{N}\approx$ 310 K, up to the maximum measured $T$ = 335 K, which could be due to the presence of known short-range magnetic ordering above $T_{N}$ in MnTe Baral et al. (2022).
@@
The structure of hexagonal $\alpha$-MnTe, with space group $P6_{3}/mmc$, is shown in Fig. 1(a). It is an A-type antiferromagnet, containing ferromagnetic planes of in-plane Mn spins that alternate spin direction along the $c$-axis. The easy axis is at 30 to the crystalline $a$-axis. Our samples are 40 nm-thick MnTe(001) thin films, grown on InP(111) via molecular beam epitaxy.
@@
We perform two-color pump-probe measurements, schematically illustrated in Fig. 1(b). We use a 1560 nm pump at 10 mW power and a 780 nm probe at 0.7 mW power on the sample; both pump and probe pulses are $\approx$100 fs in duration and have 80 MHz repetition rate. Both beams are focused through a 10x microscope objective at normal incidence, forming a 10 $\mu$m spot size. The reflected light is short-pass filtered to reject the pump, then the polarization rotation $\Delta\theta(t)$ is measured with a Wollaston prism and a balanced photodiode. Signal-to-noise is enhanced by modulating the pump with a photoelastic modulator at 42 kHz and measuring the second harmonic of the photodiode voltage output using a lock-in amplifier. Both pump and probe are linearly polarized.
@@
Figure 1: Time-resolved magneto-optical setup and observation of optical phonons in MnTe. (a) Crystal and magnetic structure of hexagonal MnTe (001). Gold and silver balls represent Mn and Te, respectively. (b) Schematic of two-color pump-probe MOKE setup with 1560 nm pump and 780 nm probe, both $\sim$100 fs in duration. (c) Probe polarization rotation $\Delta\theta(t)$ in MnTe(001) at 295 K as a function of pump polarization angle $\varphi$, and (d) associated fast Fourier transforms (FFTs).
@@
We first show short time scans (up to 13 ps) of $\Delta\theta(t)$ at room temperature (295 K) as a function of pump polarization angle $\varphi$ in Fig. 1(c). $\varphi$ is referenced to a InP(111) substrate crystalline axis; the MnTe crystalline axes should follow that of the substrate, likely with 60 crystalline twinning. After an initial sharp peak, we observe beating between two modes. Upon taking the FFTs in Fig. 1(d), we find peaks at 3.6 and 4.2 THz, whose amplitudes depend on $\varphi$. These frequencies match modes observed in previous Raman spectroscopy of MnTe Zhang et al. (2020a) that were attributed to transverse and longitudinal optical phonons, respectively. Polarization-dependent amplitudes demonstrate a non-thermal excitation mechanism, such as impulsive stimulated Raman scattering (ISRS) Yan, Gamble, and Nelson (1985) or displacive excitation of coherent phonons (DECP) Zeiger et al. (1992). Since DECP requires absorption of a photon (meaning above-bandgap excitation), ISRS does not, and our pump wavelength is below the bandgap of MnTe(1.15 eV = 1078 nm at 295 K Bossini et al. (2020)), ISRS is the most likely excitation mechanism.
Our results differ from previous time-resolved reflectivity and MOKE on MnTe Bossini et al. (2021); Deltenre et al. (2021), where only the doubly degenerate $E_{2g}$ phonon mode at 5.3 THz (also seen in Raman scattering Zhang et al. (2020a)) was observed. In these works, however, the pump wavelength was 800 nm (above the bandgap), the probe wavelength was 730 nm, and the pump-probe signal was explicitly attributed to DECP. These different excitation mechanisms and probe wavelengths might cause the differences in observed phonons.
To observe spin oscillations, we fix $\varphi=225^{\circ}$ from horizontal, where both phonon modes are excited in roughly equal magnitude, and measure $\Delta\theta(t)$ while varying the applied magnetic field $H_{app}$. To observe MO signal from out-of-plane magnetization or potential in-plane altermagnetism Gonzalez Betancourt et al. (2023), we apply out-of-plane field with a permanent magnet placed behind the sample. $\Delta\theta(t)$ scans are shown in Fig. 2(a) and associated FFTs are shown in Fig. 2(b). After the initial fast phonon oscillations, we observe slower oscillations whose amplitude depends on $H_{app}$, and do not appear when $H_{app}=0$. Oscillations of similar frequencies were observed in time-resolved transmissivity at 800 nm pump wavelength in MnTe/SrF2(111) in Ref. Zhu et al., 2023, where they were tentatively attributed to acoustic phonons 111 Note that GHz-scale acoustic phonons can be excited by laser light both above the bandgap, by the deformation potential mechanism (similar to DECP), and below the bandgap, by thermoelastic or electrostriction mechanisms (see P. Ruello and V. E. Gusev, Ultrasonics 56, 21-35 (2015) for a review.) Thus, the observation of similar-frequency acoustic phonons in both our work and Ref. Zhu et al., 2023 does not necessarily contradict the differing optical phonon measurements in our work and Ref. Bossini et al., 2021.
@@
Figure 2: Magnon and acoustic phonon oscillations at 295 K. (a) Longer scans (133 ps duration) at fixed $\varphi=225^{\circ}$ as a function of field $H_{app}$, and (b) associated FFTs. We observe a mode at 55 GHz with field-dependent amplitude, which indicates spin oscillations, as well as modes at $\sim$17 and 34 GHz that may be due to acoustic phonons. (c) Pump polarization dependence of $\Delta\theta(t)$ at fixed $H_{app}=2.2$ kG, with FFTs in (d). Oscillation amplitudes are $\varphi$-dependent, demonstrating a non-thermal origin such as phonon-induced magnons. (e) Difference and sum of $\Delta\theta(t)$ at +2.2 kG and -2.2 kG, and (f) associated FFTs. $\Delta\theta(t)$ is largely even in $H_{app}$.
@@
We fit the data from 10 ps onwards, discarding the initial phonon modes, to the sum of an exponential decay representing electron-phonon relaxation Bossini et al. (2021), two exponentially damped sinusoids representing acoustic phonons and/or magnons Gamaly and Rode (2013), and a linear background:
$\begin{split}\Delta\theta(t)=A_{exp}e^{-t/t_{0}}+A_{1}e^{-\alpha_{1}t}\sin\left(2\pi f_{1}t+\phi_{1}\right)&\\ +A_{2}e^{-\alpha_{2}t}\sin\left(2\pi f_{2}t+\phi_{2}\right)+Bt+C\end{split}$ (1)
@@
Fits are shown with the data in Fig. 2(a), except at $H_{app}$ = 0 kG, where a consistent fit could not be obtained. At this pump angle (225), the features of the data are well-captured by two damped oscillations and fitting to three did not significantly improve the fit (see the supporting information (SI)). Best fit parameters are shown in Table 1. We find that although both oscillations go to zero when $H_{app}=0$, $A_{1}$ monotonically depends on the magnitude of $H_{app}$, which suggests a magnetic origin, whereas $A_{2}$ does not.
Previous studies of magnons in MnTe are scarceSzuszkiewicz et al. (2005); Mu et al. (2019). We can estimate the $k=0$ magnon frequencies from the anisotropy, which in MnTe is largely easy-plane because optics probe the $k=0$ magnons. Since the in-plane anisotropy field $H_{Az}$ is small compared to the hard-axis (easy-plane) field $H_{Ax}$, we can use the formula for the two $k=0$ easy-plane magnon frequencies $\omega_{\alpha}$ and $\omega_{\beta}$ Rezende, Azevedo, and Rodríguez-Suárez (2019):
@@
where $\gamma=g\mu_{B}/\hbar$ is the gyromagnetic ratio and $H_{E}$ is the exchange field. For $H_{Ax}>>H_{Az}$, Eqn. 3 reduces to:
@@
Although the values of $H_{Az}$ and $H_{Ax}$ are not well-known, we can get an estimate by using the spin-flop field $H_{SF}\approx\sqrt{2H_{E}H_{Az}}$, about 2 T for MnTe/InP(111) Kriegner et al. (2017). Assuming $g=2$ for MnTe, the in-plane frequency $\omega_{\beta 0}$ evaluates to 56 GHz, consistent with our measured value. From the lack of any measurable magnetization response in MnTe with OOP field up to 6 T in Ref. Kriegner et al., 2016, we can estimate that $H_{Ax}>6$ T, which means that the OOP magnon frequency $\omega_{\alpha 0}>$ 176 GHz - therefore, features between 17 and 35 GHz are unlikely to be magnons. They may represent acoustic phonons, or a non-monotonic magneto-optical response to sample heating Yang et al. (2019). In previous studies of the easy-plane antiferromagnet NiO, the in-plane mode was observed with a linearly polarized probe through Faraday rotation, while the out-of-plane mode was observed with a circularly polarized probe through the Cotton-Mouton effect (linear birefringence) Tzschaschel et al. (2017). If similar mechanisms are active in MnTe, the higher-frequency magnon may only appear when using a circularly polarized probe, which will be the subject of future work.
@@
Following the field dependence, we measure pump polarization dependence of the spin oscillations at fixed $H_{app}$ = 2.2 kG in Fig. 2(c), with associated FFTs in 2(d). We find that the oscillation amplitudes are pump polarization-dependent, although at the current signal-to-noise level, we cannot definitively determine the symmetry. Data at $\varphi=75^{\circ}$ fits better to three oscillations, while data at $\varphi=195^{\circ}$ and 225 seems to contain only two (see SI). This suggests that multiple acoustic phonons with different pump angle dependences may be excited by the laser.
@@
In general, magnons measured with ultrafast optical pump-probe can be thermally or non-thermally generated. In thermal generation, laser heating temporarily changes the anisotropy field $H_{A}$, which reorients $H_{eff}$ and causes the spins to precess around the new $H_{eff}$ Lyalin, Cheng, and Kawakami (2021). Non-thermal magnon generation is often caused by spin-phonon coupling in magnetic semiconductors and insulators when the laser coherently excites phonons Berk et al. (2019), and was observed in MnTe in Refs. Bossini et al., 2021; Deltenre et al., 2021. In those experiments, the off-normally incident laser excited the 5.3 THz $E_{2g}$ optical phonon, causing the Te atoms to oscillate out-of-phase. This modulated the Mn-Te-Mn bond angle, in turn causing an oscillation in the weak Te-mediated superexchange $J_{3}$ which generated magnons at the same frequency. In our data, however, the 3.6 THz and 4.2 THz oscillations in $\Delta\theta(t)$ show no field dependence within experimental resolution (see the supplemental information). Therefore, we attribute them entirely to optical phonons. The polarization dependence in Fig. 2(c) is consistent with a hypothesis of acoustic phonon-induced spin oscillations at 55 GHz, driven by ISRS Yan, Gamble, and Nelson (1985).
@@
In many pump-probe experiments, magnetic and non-magnetic signals - for example, from acoustic phonons or phonon-induced spin tilting Lattery (2020); Zhang et al. (2020b) - are separated by taking the sum and difference of data acquired at opposite magnetic fields McGill et al. (2004). Therefore, in Fig. 2(e) we plot the sum and difference of $\Delta\theta(t)$ at $H_{app}=+2.2$ kG and -2.2 kG, with FFTs in 2(f). We find greater amplitude in the sum than the difference, and in fact the difference nearly cancels out after 40 ps. This means that although $\Delta\theta(t)$ is field-dependent, it does not change sign with $H_{app}$ (which can be seen as well in Fig. 2(a)). $\Delta\theta(t)$ that is even in $H_{app}$ can occur if $\Delta\theta(t)$ is due to a Voigt effect Tzschaschel et al. (2017) or optical birefringence Xu et al. (2019), which is quadratic in magnetization $m$, rather than the Kerr effect that is linear in $m$. Since the net $m$ is zero in antiferromagnets, the dominant magneto-optical effect is often these quadratic terms. Static Kerr and circular dichroism measurements of these samples showed no hysteresis Chilcote et al. , which is consistent with a lack of observed TR-MOKE signal.
After field and polarization dependence, we measure $\Delta\theta(t)$ as a function of temperature through $T_{N}$ ($\approx$ 300 K in our thin films Chilcote et al. , slightly lowered from bulk $T_{N}$ = 310 K due to strain), where long-range ordering disappears. Fig. 3(a) shows $\Delta\theta(t)$ at fixed $H_{app}$ = +2.2 kG from 295 K to 335 K. Surprisingly, oscillations at 55 GHz seem to persist at least up to 335 K, the highest temperature measured. To determine if they are still magnetic above the nominal $T_{N}$, we compare $\Delta\theta(t)$ taken at $H_{app}=$ +2.2 kG, 0 kG, and -2.2 kG at 295 K and 325 K in Fig. 3(b) and (c), respectively. Fit parameters are shown in the SI. At both temperatures, the 55 GHz mode appears only when $H_{app}\neq 0$, which indicates that it is still magnetic. Previous studies in both ferromagnets Qin et al. (2017) and antiferromagnets Biesner et al. (2022) have shown that in materials with short-range ordering above the Curie or Néel temperatures, respectively, magnetic anisotropy can persist above $T_{C}$ or $T_{N}$, and therefore collective spin oscillations can be observed at the same frequency through the transition. Since short-range magnetic order is known to exist in MnTe well above $T_{N}$ Baral et al. (2022), it is therefore plausible that MnTe continues to exhibit spin oscillations for some temperature range above $T_{N}$.
Figure 3: Measuring $\Delta\theta(t)$ through $T_{N}$. (a) $\Delta\theta(t)$ at fixed $H_{app}$ = +2.2 kG from 295 K to 335 K. 55 GHz oscillations appear at all temperatures. (b), (c) $\Delta\theta(t)$ at $H_{app}$ = +2.2 kG, 0 kG, and -2.2 kG OOP at 295 K and 325 K, respectively. At both temperatures, the 55 GHz mode appears only when $H_{app}\neq 0$, which indicates that it is magnetic.
@@
Lastly, we measure $\Delta\theta(t)$ as a function of temperature from 10 K up to 350 K in a cryostat. In this configuration, the maximum attainable $H_{app}$ was $\approx$ 500 G, which we found was not sufficient to resolve spin oscillations. Instead, we focus on the optical phonon oscillations, showing $\Delta\theta(t)$ traces in Fig. 4(a) and FFTs in 4(b). (Full time traces showing no spin oscillations are shown in the SI.)
@@
To analyze the temperature dependence of the phonons, we fit the background, then subtract it off and fit the residual oscillations to two damped sinusoids (see SI). We find that this two-step process is more robust than fitting the background and oscillations simultaneously. At low temperatures, modeling the background by a sum of decaying exponentials does not yield a satisfactory fit. Therefore, we adapt a phenomenological model developed to fit pump-probe reflectivity in Ref. Djordjevic et al., 2006: